### v1

# Higher-Rank Irreducible Cartesian Tensors for Equivariant Message Passing

Abstract
The ability to perform fast and accurate atomistic simulations is crucial for advancing the chemical sciences. By learning from high-quality data, machine-learned interatomic potentials achieve accuracy on par with ab initio and first-principles methods at a fraction of their computational cost. The success of machine-learned interatomic potentials arises from integrating inductive biases such as equivariance to group actions on an atomic system, e.g., equivariance to rotations and reflections. In particular, the field has notably advanced with the emergence of equivariant message-passing architectures. Most of these models represent an atomic system using spherical tensors, tensor products of which require complicated numerical coefficients and can be computationally demanding. This work introduces higher-rank irreducible Cartesian tensors as an alternative to spherical tensors, addressing the above limitations. We integrate irreducible Cartesian tensor products into message-passing neural networks and prove the equivariance of the resulting layers. Through empirical evaluations on various benchmark data sets, we consistently observe on-par or better performance than that of state-of-the-art spherical models.
1 Introduction
The ability to perform sufficiently fast and accurate atomistic simulations for large molecular and material systems holds the potential to revolutionize molecular and materials science [1, 2, 3, 4, 5, 6, 7, 8]. Conventionally, computational chemistry and materials science rely on ab initio or first-principles approaches—e.g., coupled cluster [9, 10, 11] or density functional theory (DFT) [12, 13], respectively—that are accurate but computationally demanding, thus limiting the accessible simulation time and system sizes. However, the ability to generate high-quality, first-principles-based data sets has prompted the development of machine-learned interatomic potentials (MLIPs). These potentials enable atomistic simulations with accuracy that is on par with the reference first-principles method but at a fraction of the computational cost. Message-passing neural networks (MPNNs) [14, 15, 16, 17, 18, 19, 20, 21] have been employed in chemical and materials sciences, including the development of MLIPs, due to their efficient processing of the graph representation of the atomic system [22, 23, 24, 25, 26]. Achieving the desired MLIPs’ performance, however, requires the inclusion of inductive biases, like the invariance of total energy to translations, reflections, and rotations in the three-dimensional space, and an effective encoding of the atomic system into a learnable representation [27]. Designing equivariant MPNNs [28, 29, 30, 31, 32, 25, 33, 34, 35, 36, 26], which preserve the directional information of the local atomic environment, has been one of the most active research directions of the last years. Previous work often achieves equivariance by representing an atomic system as a graph where node features are expressed in spherical harmonics basis and designing ad-hoc message-passing layers [25, 32, 33, 35, 34, 36].
MLIPs based on spherical harmonics [25, 33, 34], which are the basis for irreducible representations of the three-dimensional rotation group, often demonstrate a better performance compared to those that use lower-rank Cartesian representations (scalars and vectors) [28, 37]. Spherical tensors, however, require the definition of a particular rotational axis, often chosen as the $z$-axis, resulting in inherent bias [38]. Furthermore, coupling spherical tensors via tensor products, involved in designing equivariant convolutions and constructing many-body features [39, 40, 33], requires the definition of complicated numerical coefficients, e.g., Wigner $3\mbox{-}j$ symbols defined in terms of the Clebsch–Gordan coefficients [41], and can be computationally demanding. In contrast, irreducible Cartesian tensors have no preferential directions; their tensor products are simpler and have a better computational complexity—up to a certain tensor rank—than the tensor products of spherical tensors [38, 42, 43, 44, 45]. Recent work [30] improved results of Cartesian MPNNs by decomposing rank-two tensors into irreducible representations of the three-dimensional rotation group, i.e., into representations of dimension 1 (trace), 3 (anti-symmetric part), and 5 (traceless symmetric part) [46].
Contributions. In this work, we demonstrate that operating with irreducible Cartesian tensors leads to on-par or, sometimes, even better performance than that of spherical counterparts. Particularly, this work goes beyond scalars, vectors, and rank-two tensors and explores the integration of higher-rank irreducible Cartesian tensors and their products into equivariant MPNNs: i) We demonstrate how irreducible Cartesian tensors that are symmetric and traceless can be constructed from a unit vector and a product of two irreducible Cartesian tensors; ii) We prove that the resulting tensors are equivariant under the action of the three-dimensional rotation and reflection group; iii) We demonstrate that higher-rank irreducible Cartesian tensors can be used to design cost-efficient—up a certain tensor rank—and accurate equivariant models of many-body interactions; iv) We conduct different experiments to assess the effectiveness of equivariant message passing based on irreducible Cartesian tensors, and achieve state-of-the-art performance on benchmark data sets such as rMD17 [47], 3BPA [48], and acetylacetone [32]. We hope our contributions will offer new insights into the use of Cartesian tensors for constructing accurate and cost-efficient MLIPs and beyond.
2 Background
Group representations and equivariance. A group $(G,\cdot)$ is defined by a set of elements $G$ and a group product $\cdot$. A representation $D$ of a group is a function from $G$ to square matrices such that $D[g]D[g^{\prime}]=D[g\cdot g^{\prime}],\,\forall\,g,g^{\prime}\in G$. This representation defines an action of $G$ to any vector space $\mathcal{X}$ (of the same dimension as the dimension of square matrices) through the matrix-vector multiplication, i.e., $(g,\mathbf{x})\mapsto D_{\mathcal{X}}[g]\mathbf{x}$ for $\forall\,g\in G$ and $\forall\,\mathbf{x}\in\mathcal{X}$. For vector spaces $\mathcal{X}$ and $\mathcal{Y}$, a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ is called equivariant to the action of a group $G$ to $\mathcal{X}$ and $\mathcal{Y}$ iff
$f(D_{\mathcal{X}}[g]\mathbf{x})=D_{\mathcal{Y}}[g]f(\mathbf{x})\,,\forall\,g\in G.$
Invariance can be seen as a special type of equivariance with $D_{\mathcal{Y}}[g]$ being the identity for all $g$. An important class of equivariant models focuses on equivariance to the action of the Euclidean group $\mathrm{E}(3)$, comprising translations and the orthogonal group $\mathrm{O}(3)$, i.e., rotation group $\mathrm{SO}(3)$ and reflections, in $\mathbb{R}^{3}$. MLIPs based on equivariant models usually focus on equivariance to the action of the orthogonal group $\mathrm{O}(3)$ and are invariant to translations.
Cartesian tensors. A vector $\mathbf{x}\in\mathbb{R}^{3}$ transforms under the action of the rotation group $\mathrm{SO}(3)$ as $\mathbf{x}^{\prime}=R\mathbf{x}$, i.e., each component of it transforms as $\left(\mathbf{x}\right)_{i}^{\prime}=\sum_{j}R_{ij}\left(\mathbf{x}\right)_{j}$. Here, $R=D_{\mathcal{X}}\left[g\right]\in\mathbb{R}^{3\times 3}$ denotes the rotation matrix representation of $g\in\mathrm{SO}(3)$. Cartesian tensors of rank $n$ are described by $3^{n}$ numbers and generalize the concept of vectors. A rank-$n$ Cartesian tensor $\mathbf{T}$ can be viewed as a multidimensional array with $n$ indices, i.e., $\left(\mathbf{T}\right)_{i_{1}i_{2}\cdots i_{n}}$ with $i_{k}\in\{1,2,3\}$ for $\forall\,k\in\{1,\cdots,n\}$. Furthermore, each index of $\left(\mathbf{T}\right)_{i_{1}i_{2}\cdots i_{n}}$ transforms independently as a vector under rotation. For example, a rank-two tensor transforms under rotation as $\mathbf{T}^{\prime}=R\mathbf{T}R^{\top}$, i.e., each component of it transforms as $\left(\mathbf{T}\right)_{i_{1}i_{2}}^{\prime}=\sum_{j_{1}j_{2}}R_{i_{1}j_{1}}R_{i_{2}j_{2}}\left(\mathbf{T}\right)_{j_{1}j_{2}}$. For Cartesian tensors, one defines an $r$-fold tensor contraction and an outer product, i.e., $\left(\mathbf{T}\right)_{j_{1}\cdots j_{s}}=\sum_{i_{1},\cdots,i_{r}}\left(\mathbf{U}\right)_{i_{1}\cdots i_{r}j_{1}\cdots j_{s}}\left(\mathbf{S}\right)_{i_{1}\cdots i_{r}}$ and $\left(\mathbf{T}\right)_{i_{1}\cdots i_{r}j_{1}\cdots j_{s}}=\left(\mathbf{U}\right)_{i_{1}\cdots i_{r}}\left(\mathbf{S}\right)_{j_{1}\cdots j_{s}}$, respectively. Cartesian tensors are generally reducible and can, thus, be decomposed into smaller representations that transform independently within their linear subspaces under rotation. For example, a rank-two reducible Cartesian tensor contains representations of dimension 1 (trace), 3 (anti-symmetric part), and 5 (traceless symmetric part): $\sum_{i_{1}}(\mathbf{T})_{i_{1}i_{1}}$, $(\mathbf{T})_{i_{1}i_{2}}-(\mathbf{T})_{i_{2}i_{1}}$, and $(\mathbf{T})_{i_{1}i_{2}}+(\mathbf{T})_{i_{2}i_{1}}-2/3\sum_{i_{1}}(\mathbf{T})_{i_{1}i_{1}}$, respectively. Under rotation, the symmetric traceless part remains within its irreducible subspace, with a similar behavior for other irreducible components of $(\mathbf{T})_{i_{1}i_{2}}$. Furthermore, for a reducible Cartesian tensor of rank $n$, an action of the rotation group $\mathrm{SO}(3)$ can be represented with a $3^{n}\times 3^{n}$-dimensional rotation matrix. This rotation matrix is also reducible, meaning that an appropriate change of basis can block diagonalize it into smaller subsets, the irreducible representations of the rotation group $\mathrm{SO}(3)$.
Spherical tensors. Spherical harmonics (or spherical tensors) $Y_{m}^{l}$ are functions from the points on a sphere to complex or real numbers, with degree $l\geq 0$ and components $-l\leq m\leq l$. Collecting all components for a given $l$ we obtain a $(2l+1)$-dimensional object $\mathbf{Y}^{l}=(Y_{-l}^{l},\dots,Y_{l-1}^{l},Y_{l}^{l})$. Spherical tensors transform under rotation as $Y_{m}^{l}(R\hat{\mathbf{x}})=\sum_{m^{\prime}}(\mathbf{D})_{mm^{\prime}}^{l}Y_{m^{\prime}}^{l}(\hat{\mathbf{x}})$, where $(\mathbf{D})_{mm^{\prime}}^{l}$ is the irreducible representation of $\mathrm{SO}(3)$—the Wigner $D$-matrix—and $R$ denotes the rotation matrix. Spherical tensors are irreducible and can be combined using the Clebsch–Gordan tensor product
$\left(Y_{m_{1}}^{l_{1}}\otimes_{\mathrm{CG}}Y_{m_{2}}^{l_{2}}\right)_{m_{3}}^{l_{3}}=\sum\nolimits_{m_{1}=-l_{1}}^{l_{1}}\sum\nolimits_{m_{2}=-l_{2}}^{l_{2}}C_{l_{1}m_{1},l_{2}m_{2}}^{l_{3}m_{3}}Y_{m_{1}}^{l_{1}}Y_{m_{2}}^{l_{2}},$
with Clebsch–Gordan coefficients $C_{l_{1}m_{1},l_{2}m_{2}}^{l_{3}m_{3}}$ and $l_{3}\in\left\{\lvert l_{1}-l_{2}\rvert,\cdots,l_{1}+l_{2}\right\}$. Furthermore, any reducible Cartesian tensor of rank $n$ can be written in spherical harmonics as $\left(\hat{\mathbf{x}}\right)_{i_{1}}\left(\hat{\mathbf{x}}\right)_{i_{2}}\cdots\left(\hat{\mathbf{x}}\right)_{i_{n}}=\sum_{l=0}^{n}\sum_{m=-l}^{l}X_{i_{1}i_{2}\cdots i_{n}}^{lm}Y_{m}^{l}$, with coefficients $X_{i_{1}i_{2}\cdots i_{n}}^{lm}$ defined elsewhere [49].
3 Related work
Higher-body-order local descriptors. Recent advances in MLIPs have been influenced by moment tensor potentials (MTP) [50] and atomic cluster expansion (ACE) [40]. These approaches enable systematic construction of higher-body-order polynomial basis functions, encompassing representations like atom-centered symmetry functions (ACSFs) [51, 52], smooth overlap of atomic positions (SOAP) [53, 54], Gaussian moments [55, 56], and embedded atom/multi-scale embedded atom method (EAM/MEAM) potentials [57, 58]. Furthermore, a reducible Cartesian tensor can be represented as a linear combination of irreducible spherical counterparts, and vice versa [40, 49]. More general expressions, including tensor contractions, have also been provided, hinting at the relationship between Cartesian and spherical models [59, 60, 61]. Despite the success of MTP and ACE, defining smaller cutoff radii and rigid architecture can result in limited accuracy compared to MPNNs.
Equivariant message-passing potentials. Equivariant MPNNs [39, 62, 28, 63, 64, 25, 37, 32, 33, 29, 35, 34, 30, 36, 31] often outperform more traditional invariant counterparts [65, 66, 67, 68, 69, 70]. While many invariant and equivariant MPNNs rely on two-body features within a single message-passing layer, there is a growing interest in incorporating higher-body-order features, such as angles and dihedrals, to model many-body interactions in atomic systems [69, 70, 67]. Note that equivariant models with two-body features in a single message-passing layer build many-body ones through numerous message-passing iterations, increasing the receptive field and, thus, computational cost. Refs. 67, 69 recognized the importance of higher-body-order features but faced challenges due to explicit summation over triplets or quadruplets. In contrast, MACE advances the current state-of-the-art by combining ACE and equivariant message passing, introducing cost-efficient many-body message-passing potentials [33]. With just two message-passing layers, MACE yields accurate potentials for interacting many-body systems [71].
Beyond Clebsch–Gordan tensor product. Despite the success of equivariant MPNNs based on spherical tensors, the high computational cost of the Clebsch–Gordan tensor product limits their computational efficiency [72, 28, 63, 73, 33, 64, 34, 35, 30, 74]. Thus, current research focuses on alternatives to the Clebsch–Gordan tensor product, which has a $\mathcal{O}(L^{6})$ complexity for tensors up to degree $L$. Recently, the relation of Clebsch–Gordan coefficients to the integral of products of three spherical harmonics, known as the Gaunt coefficients [41], has been exploited to reduce the computational cost of the tensor product of spherical tensors to $\mathcal{O}(L^{3})$ [36]. Alternatively, Cartesian tensors and their products can be used to construct message-passing layers with two- and many-body features. Recently, a competitive performance to spherical models for a Cartesian message-passing architecture based on decomposing rank-two tensors into their irreducible representations has been demonstrated [30]. In this case, all operations are performed on tensors or their decomposition, which has advantageous computational complexity compared to the Clebsch–Gordan tensor product. Furthermore, recent work employs higher-rank reducible Cartesian tensors to design many-body message-passing potentials but does not achieve the state-of-the-art performance of spherical counterparts [31]. This performance gap may result from mixing different irreducible representations during massage-passing based on reducible Cartesian tensors and their products.
4 Methods
We define an atomic configuration $\mathcal{S}=\{\mathbf{r}_{u},Z_{u}\}_{u=1}^{N_{\mathrm{at}}}$, where $\mathbf{r}_{u}\in\mathbb{R}^{3}$ denotes Cartesian coordinates and $Z_{u}\in\mathbb{N}$ represents the atomic number of atom $u$, with a total of $N_{\mathrm{at}}$ atoms. Our focus lies on message-passing MLIPs, parameterized by $\boldsymbol{\theta}$, that learn a mapping from a configuration $\mathcal{S}$ to the total energy $E$, i.e., $f_{\boldsymbol{\theta}}:\mathcal{S}\mapsto E\in\mathbb{R}$. Thus, we represent molecular and material systems as graphs in a three-dimensional Euclidean space. An edge $\{u,v\}$ exists if atoms $u$ and $v$ are within a cutoff distance $r_{\mathrm{c}}$, i.e., $\lVert\mathbf{r}_{u}-\mathbf{r}_{v}\rVert_{2}\leq r_{\mathrm{c}}$. For more details on MPNNs, see Appendix A. The total energy of an atomic configuration is defined by the sum of individual atomic energy contributions [51], i.e., $E=\sum_{u=1}^{N_{\mathrm{at}}}E_{u}$. Atomic forces are computed as negative gradients of the total energy with respect to atomic coordinates, i.e., $\mathbf{F}_{u}=-\nabla_{\mathbf{r}_{u}}E$.
4.1 Irreducible Cartesian tensor product
Figure 1: Schematic illustration of (a) the construction of an irreducible Cartesian tensor for a local atomic environment and (b) the tensor product of two irreducible Cartesian tensors of rank $l_{1}$ and $l_{2}$. The construction of an irreducible Cartesian tensor from a unit vector $\hat{\mathbf{r}}$ is defined in Eq. (1). In this work, we use tensors with the same rank $n$ and weight $l$, i.e., $n=l$, avoiding the need for embedding tensors with $l<n$ in a higher-dimensional tensor space. Therefore, we use $l$ to identify the rank and the weight of an irreducible Cartesian tensor throughout this work. The tensor product is defined in Eqs. (2) and (3), resulting in a new tensor $\mathbf{T}_{l_{3}}=(\mathbf{T}_{l_{1}}\otimes_{\mathrm{Cart}}\mathbf{T}_{l_{2}})_{l_{3}}$ of rank $l_{3}=\{\lvert l_{1}-l_{2}\rvert,\cdots,l_{1}+l_{2}\}$. Transparent boxes denote the linearly dependent elements of symmetric and traceless tensors. The tensor product yields even and odd parity tensors $\mathbf{T}_{l_{3}}$, given by $l_{1}+l_{2}-l_{3}$.
In the following, we explore irreducible Cartesian tensors and their products based on the three-dimensional vector space and the three-dimensional orthogonal group $\mathrm{O}(3)$, which comprises rotations and reflections. The respective tensors and tensor products are schematically illustrated in Fig. 1 (a) and (b), respectively, and are further employed in constructing MPNNs for atomic systems equivariant under actions of the orthogonal group. An irreducible Cartesian tensor $\mathbf{T}_{n}\in(\mathbb{R}^{3})^{\otimes n}$ of rank $n$ and weight $l\leq n$ (related to the degree $l$ of spherical tensors) can be represented by a tensor with $3^{n}$ components in a three-dimensional vector space. These $3^{n}$ components form a basis for a $(2l+1)$-dimensional irreducible representation of the three-dimensional rotation group $\mathrm{SO}(3)$ [75, 76]. Thus, only $12l+$ 1 of the $3^{n}$ components are independent.
The rotation in the space of all tensors of rank $n$ is induced through the $n$-fold outer product of a rotation matrix $R\in\mathbb{R}^{3\times 3}$, i.e., $R^{\otimes n}=R\otimes\cdots\otimes R$. The obtained representation of the rotation group $\mathrm{SO}(3)$ is reducible for all $n$ except $n=\{0,1\}$. Reducing a tensor of rank $n$ yields a unique irreducible tensor with the same weight and rank ($n=l$), which is characterized by being symmetric, i.e., $(\mathbf{T}_{n})_{\cdots i\cdots j\cdots}=(\mathbf{T}_{n})_{\cdots j\cdots i\cdots}\ ,\forall\ i\neq j\in\{i_{1}\cdots,i_{n}\}$, and traceless, i.e., $\sum_{i}(\mathbf{T}_{n})_{\cdots i\cdots i\cdots}=0\ ,\forall\,i\in\{i_{1},\cdots,i_{n}\}$. An irreducible tensor of rank $n$ and weight $l$ with $l<n$ can be viewed as a $l$-rank tensor embedded in the $n$-rank tensor space, e.g., by computing an outer product with the identity matrix. However, the embedding is often not unique. Thus, we construct tensors with the same weight and rank ($n=l$) in the following.
Irreducible Cartesian tensors from unit vectors. Hereafter, we denote the rank and the weight of irreducible Cartesian tensors by $l$ to distinguish them from reducible counterparts. An irreducible Cartesian tensor of an arbitrary rank $l$ can be constructed from a unit vector $\hat{\mathbf{r}}$ in the form [45]
$\mathbf{T}_{l}(\hat{\mathbf{r}})=C\sum\nolimits_{m=0}^{\lfloor l/2\rfloor}(-1)^{m}\frac{(2l-2m-1)!!}{(2l-1)!!}\big{\{}\hat{\mathbf{r}}^{\otimes(l-2m)}\otimes\mathbf{I}^{\otimes m}\big{\}},$ (1)
resulting in a symmetric and traceless tensor of rank $l$. Here, $\mathbf{I}$ denotes the $3\times 3$ identity matrix, $\hat{\mathbf{r}}^{\otimes(l-2m)}=\hat{\mathbf{r}}\otimes\cdots\otimes\hat{\mathbf{r}}$ and $\mathbf{I}^{\otimes m}=\mathbf{I}\otimes\cdots\otimes\mathbf{I}$ are the corresponding $(l-2m)$- and $m$-fold outer products. The curly brackets indicate the summation over all permutations of the $l$ unsymmetrized indices [45], i.e., $\{\mathbf{T}_{l}\}_{i_{1}\cdots i_{l}}=\sum_{\pi\in S_{l}}(\mathbf{T}_{l})_{i_{\pi(1)}\cdots i_{\pi(l)}}$, with $S_{l}$ being the corresponding set of permutations. The expression in Eq. (1) involves three distinct outer products $(\hat{\mathbf{r}}\otimes\hat{\mathbf{r}})_{i_{1}i_{2}}=\hat{r}_{i_{1}}\hat{r}_{i_{2}}$, $(\mathbf{I}\otimes\mathbf{I})_{i_{1}i_{2}i_{3}i_{4}}=\delta_{i_{1}i_{2}}\delta_{i_{3}i_{4}}$, and $(\hat{\mathbf{r}}\otimes\mathbf{I})_{i_{1}i_{2}i_{3}}=\hat{r}_{i_{1}}\delta_{i_{2}i_{3}}$, where $\delta_{i_{2}i_{3}}$ denotes the Kronecker delta. The normalization constant $C=\left(2l-1\right)!!/l!$ is chosen such that an $l$-fold contraction of $\mathbf{T}_{l}$ with the unit vector $\hat{\mathbf{r}}$ yields unity. An example of an irreducible Cartesian tensor with rank $l=3$ is $(\mathbf{T}_{l=3})_{i_{1}i_{2}i_{3}}=\frac{5}{2}\big{(}\hat{r}_{i_{1}}\hat{r}_{i_{2}}\hat{r}_{i_{3}}-\frac{1}{5}(\hat{r}_{i_{1}}\delta_{i_{2}i_{3}}+\hat{r}_{i_{2}}\delta_{i_{3}i_{1}}+\hat{r}_{i_{3}}\delta_{i_{1}i_{2}})\big{)}$.
Irreducible Cartesian tensor product. The following defines the product of two irreducible Cartesian tensors $\mathbf{x}_{l_{1}}\in(\mathbb{R}^{3})^{\otimes l_{1}}$ and $\mathbf{y}_{l_{2}}\in(\mathbb{R}^{3})^{\otimes l_{2}}$, yielding an irreducible Cartesian tensor of rank $l_{3}$, i.e., $\mathbf{z}_{l_{3}}=(\mathbf{x}_{l_{1}}\otimes_{\mathrm{Cart}}\mathbf{y}_{l_{2}})_{l_{3}}\in(\mathbb{R}^{3})^{\otimes l_{3}}\,\forall\,l_{3}\in\{\lvert l_{1}-l_{2}\rvert,\cdots,l_{1}+l_{2}\}$, that is symmetric and traceless. The irreducible Cartesian tensor product is crucial for designing equivariant MPNNs and is used for equivariant convolutions and constructing many-body features in Section 4.2. For an even $l_{1}+l_{2}-l_{3}=2k$, the general form of an irreducible Cartesian tensor of rank $l_{3}$ reads [45]
$\begin{split}(\mathbf{x}_{l_{1}}&\otimes_{\mathrm{Cart}}\mathbf{y}_{l_{2}})_{l_{3}}=\\ &C_{l_{1}l_{2}l_{3}}\sum\nolimits_{m=0}^{\mathrm{min}(l_{1},l_{2})-k}(-1)^{m}2^{m}\frac{(2l_{3}-2m-1)!!}{(2l_{3}-1)!!}\big{\{}\left(\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)\otimes\mathbf{I}^{\otimes m}\big{\}},\end{split}$ (2)
where $\left(\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)=\sum_{i_{1},\cdots,i_{k+m}}(\mathbf{x}_{l_{1}})_{i_{1}\cdots i_{k+m}}(\mathbf{y}_{l_{2}})_{i_{1}\cdots i_{k+m}}$ denotes an $(k+m)$-fold tensor contraction, which results in a tensor of rank $l_{1}+l_{2}-2(k+m)$. For simplicity, we skip the uncontracted indices in the above definition. For example, for $l_{1}=4$ and $l_{2}=3$ and a three-fold tensor contraction we obtain $\left(\mathbf{x}_{l_{1}}\cdot 3\cdot\mathbf{y}_{l_{2}}\right)_{i_{4}}=\sum_{i_{1},i_{2},i_{3}}(\mathbf{x}_{l_{1}})_{i_{1}i_{2}i_{3}i_{4}}(\mathbf{y}_{l_{2}})_{i_{1}i_{2}i_{3}}$, i.e., the corresponding tensors are contracted along $i_{1}$, $i_{2}$, and $i_{3}$. Note that the final result is independent of the index permutation, as the contracted tensors are symmetric. For an odd $l_{1}+l_{2}-l_{3}=2k+1$, we define [45]
$\begin{split}(\mathbf{x}_{l_{1}}&\otimes_{\mathrm{Cart}}\mathbf{y}_{l_{2}})_{l_{3}}=\\ &D_{l_{1}l_{2}l_{3}}\sum\nolimits_{m=0}^{\mathrm{min}(l_{1},l_{2})-k-1}(-1)^{m}2^{m}\frac{(2l_{3}-2m-1)!!}{(2l_{3}-1)!!}\big{\{}\left(\boldsymbol{\varepsilon}:\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)\otimes\mathbf{I}^{\otimes m}\big{\}},\end{split}$ (3)
with $\boldsymbol{\varepsilon}$ denoting the Levi-Civita symbol ($\varepsilon_{i_{1}i_{2}i_{3}}=-\varepsilon_{i_{3}i_{2}i_{1}}$ and $\varepsilon_{i_{1}i_{1}i_{3}}=0$). The double contraction with the Levi-Civita symbol reads $(\boldsymbol{\varepsilon}:\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}})_{i_{1}}=\sum_{i_{2},i_{3}}\varepsilon_{i_{1}i_{2}i_{3}}\left(\mathbf{x}_{l_{1}}\cdot(k+m)\cdot\mathbf{y}_{l_{2}}\right)_{i_{2}i_{3}}$, and yields a tensor of rank $l_{1}+l_{2}-2(k+m)-1$. Details on the normalization constants $C_{l_{1}l_{2}l_{3}}$ and $D_{l_{1}l_{2}l_{3}}$ are provided in Appendix B.1.
4.2 Equivariant message-passing based on irreducible Cartesian tensors
The following section introduces the basic operations for constructing equivariant MPNNs based on irreducible Cartesian tensors. Using their irreducible tensor products, we also demonstrate how to build equivariant two- and many-body features, crucial for modeling many-body interactions in molecular and materials systems. Particularly, we focus on MLIPs based on equivariant MPNNs and extend the state-of-the-art MACE architecture [33] to the Cartesian basis. Following the MACE architecture, we use only tensors and tensor products of even parity. We split vectors $\mathbf{r}_{uv}=\mathbf{r}_{u}-\mathbf{r}_{v}\in\mathbb{R}^{3}$ from atom $u$ to atom $v$, schematically shown in Fig. 1, into their radial and angular components (unit vectors), i.e., $r_{uv}=\lVert\mathbf{r}_{uv}\rVert_{2}\in\mathbb{R}$ and $\hat{\mathbf{r}}_{uv}=\mathbf{r}_{uv}/r_{uv}\in\mathbb{R}^{3}$, respectively. In the $t$-th message-passing layer, edges $\{u,v\}$ are embedded using a fully connected neural network $R_{kl_{1}l_{2}l_{3}}^{(t)}:\mathbb{R}\rightarrow\mathbb{R}$ with $k$ output feature channels. The radial function $R_{kl_{1}l_{2}l_{3}}^{(t)}$ takes as an input radial distances $r_{uv}$, which are embedded through Bessel functions and multiplied by a smooth polynomial function [70]. Finally, we use irreducible Cartesian tensors $\mathbf{T}_{l}(\hat{\mathbf{r}})$, similar to spherical tensors $\mathbf{Y}^{l}(\hat{\mathbf{r}})$, to embed unit vectors into the tensor space of maximal rank $l_{\mathrm{max}}$.
Equivariant convolutions and two-body features. Rotation equivariance in MPNNs is typically achieved by constraining convolution filters to be the products between learnable radial functions and spherical tensors, i.e., $R_{kl_{1}l_{2}l_{3}}^{(t)}(r_{uv})Y_{m_{1}}^{l_{1}}(\hat{\mathbf{r}}_{uv})$. The two-body features $A_{ukl_{3}m_{3}}^{(t)}$ are further obtained through the tensor product—the point-wise convolution [39]—between the respective filters and neighbors’ equivariant features $h_{ukl_{2}m_{2}}^{(t)}$. The permutational invariance is enforced by pooling over the neighbors $v\in\mathcal{N}(u)$. Here, we use irreducible Cartesian tensors, with the rotation-equivariant filters given by $R_{kl_{1}l_{2}l_{3}}^{(t)}(r_{uv})\big{(}\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}$. Thus, two-body features $\big{(}\mathbf{A}_{ukl_{3}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{3}}}$ are obtained using the irreducible Cartesian tensor product and are represented by rank-$l_{3}$ irreducible Cartesian tensors. The Cartesian two-body features are defined by
$\big{(}\mathbf{A}_{ukl_{3}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{3}}}=\sum\nolimits_{v\in\mathcal{N}(u)}\Big{(}R_{kl_{1}l_{2}l_{3}}^{(t)}(r_{uv})\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\otimes_{\mathrm{Cart}}\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{kk^{\prime}l_{2}}^{(t)}\mathbf{h}_{vk^{\prime}l_{2}}^{(t)}\Big{)}_{i_{1}i_{2}\cdots i_{l_{3}}},$ (4)
where $d_{t}$ represents the number of feature channels in the node embeddings $\mathbf{h}_{vk^{\prime}l_{2}}^{(t)}$ of the $t$-th message-passing layer. In the first message-passing layer, node embeddings are initialized as learnable weights $W_{kZ_{v}}$ that are invariant to actions of the orthogonal group, i.e., are scalars or tensors of rank $l_{2}=0$, and embed the atom type $Z_{v}$. Thus, constructing equivariant two-body features simplifies to
$\big{(}\mathbf{A}_{ukl_{1}}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}=\sum\nolimits_{v\in\mathcal{N}(u)}R_{kl_{1}}^{(1)}(r_{uv})\big{(}\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}W_{kZ_{v}}.$ (5)
Equivariant many-body features. The importance of many-body terms arises from the fact that the interaction between atoms changes when additional atoms are present; see Appendix A. Furthermore, many-body terms are often required to ensure the transferability of interatomic potentials [77]. Here, we construct $(\nu+1)$-body equivariant features from $\big{(}\mathbf{A}_{ukl_{\xi}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{\xi}}}$ obtained using Eqs. (4) or (5). The $\nu$-fold Cartesian tensor product, which yields $(\nu+1)$-body features represented by an irreducible Cartesian tensor of rank $L$, reads
$\big{(}\mathbf{B}_{u\eta_{\nu}kL}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}}=\big{(}\underbrace{\tilde{\mathbf{A}}_{ukl_{1}}^{(t)}\otimes_{\mathrm{Cart}}\cdots\otimes_{\mathrm{Cart}}\tilde{\mathbf{A}}_{ukl_{\nu}}^{(t)}}_{\nu\text{-fold}}\big{)}_{i_{1}i_{2}\cdots i_{L}},$ (6)
where $\eta_{\nu}$ counts all possible $\nu$-fold products of $\{l_{1},\cdots,l_{\nu}\}$-rank tensors, yielding rank-$L$ irreducible Cartesian tensors, and $\tilde{\mathbf{A}}_{ukl_{\xi}}^{(t)}=\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{kk^{\prime}l_{\xi}}^{(t)}\mathbf{A}_{uk^{\prime}l_{\xi}}^{(t)}$ with $d_{t}$ feature channels.
[…]
Runtime considerations. When choosing an architecture to implement MLIPs, the runtime per force and energy evaluation is crucial. The asymptotic time complexity as a function of rank $L$ is $\mathcal{O}\left(9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$ for irreducible Cartesian tensors and $\mathcal{O}\left(L^{6}\right)$ for spherical ones; see Appendix B.3 for more details. Thus, spherical tensors are more computationally efficient for $L\rightarrow\infty$. However, state-of-the-art models and physical properties typically require $L\leq 4$ [33, 78], making sub-leading terms and implementation-dependent amplitudes crucial. Based on our analysis, for $L\leq 4$, we can expect models based on irreducible Cartesian tensors to be more computationally and memory efficient than their spherical counterparts. Nonetheless, it is essential to consider the inference times for a fair comparison, which we provide in Section 5.
Equivariance of message-passing layers. We conclude this section by giving a theoretical result that ensures the equivariance of message-passing layers based on irreducible Cartesian tensors and their irreducible tensor products to actions of the orthogonal group. The proof is provided in Appendix C.
Proposition 4.1.
The message-passing layers based on irreducible Cartesian tensors and their irreducible tensor products are equivariant to actions of the orthogonal group.
5 Experiments and results
This section presents the results for the three benchmark data sets: rMD17, 3BPA, and acetylacetone. We describe data sets and training details in Appendix D.1 and Appendix D.2, respectively.
Table 1: Energy (E) and force (F) mean absolute errors (MAEs) for the rMD17 data set. E- and F-MAE are provided in meV and meV/Å, respectively. Results are presented for models trained using $N_{\text{train}}=950$ and $N_{\text{train}}=50$ configurations randomly drawn from the data set. Further 50 configurations are used for early stopping. All values are obtained by averaging over five independent runs, with the standard deviation provided if available. Best performances, considering the standard deviation, are highlighted in bold.
$N_{\text{train}}=950$ $N_{\text{train}}=50$
[…]
Aspirin E 2.27 $\pm$ 0.11 2.4 2.2 2.3 2.3 19.5 17.0 14.84 $\pm$ 0.98
F 6.67 $\pm$ 0.19 8.9 $\pm$ 0.1 6.6 7.3 8.2 52.0 43.9 40.19 $\pm$ 0.95
Azobenzene E 1.20 $\pm$ 0.01 0.7 1.2 1.2 0.7 6.0 5.4 5.47 $\pm$ 0.63
F 2.92 $\pm$ 0.03 3.1 3.0 2.6 2.9 20.0 17.7 17.25 $\pm$ 0.53
Benzene E 0.26 $\pm$ 0.00 0.02 0.4 0.3 0.04 0.6 0.7 0.38 $\pm$ 0.02
F 0.34 $\pm$ 0.02 0.3 0.3 0.2 0.3 2.9 2.7 2.45 $\pm$ 0.13
Ethanol E 0.43 $\pm$ 0.02 0.5 0.4 0.4 0.4 8.7 6.7 6.15 $\pm$ 0.26
F 2.63 $\pm$ 0.10 3.5 2.1 2.1 2.8 40.2 32.6 29.53 $\pm$ 1.14
Malonaldehyde E 0.82 $\pm$ 0.03 0.8 0.8 0.6 0.8 12.7 10.0 9.72 $\pm$ 0.42
F 4.96 $\pm$ 0.21 5.4 4.1 3.6 5.1 52.5 43.3 42.88 $\pm$ 3.08
Naphthalene E 0.56 $\pm$ 0.00 0.2 0.5 0.2 0.9 2.1 2.1 2.06 $\pm$ 0.10
F 1.45 $\pm$ 0.05 1.6 1.6 0.9 1.3 10.0 9.2 9.43 $\pm$ 0.46
Paracetamol E 1.44 $\pm$ 0.03 1.3 1.3 1.5 1.4 14.3 9.7 8.94 $\pm$ 0.66
F 4.89 $\pm$ 0.11 5.9 $\pm$ 0.1 4.8 4.9 5.9 39.7 31.5 30.13 $\pm$ 1.51
Salicylic acid E 0.97 $\pm$ 0.01 0.8 0.9 0.9 0.7 8.0 6.5 5.95 $\pm$ 0.43
F 3.66 $\pm$ 0.06 4.6 $\pm$ 0.1 3.1 2.9 4.0 35.9 28.4 27.78 $\pm$ 1.93
Toluene E 0.46 $\pm$ 0.00 0.3 0.5 0.4 0.3 3.3 3.1 2.45 $\pm$ 0.13
F 1.61 $\pm$ 0.02 1.7 1.5 1.8 1.6 15.1 12.1 11.24 $\pm$ 0.55
Uracil E 0.57 $\pm$ 0.01 0.4 0.5 0.6 0.4 7.3 4.4 4.66 $\pm$ 0.16
F 2.64 $\pm$ 0.08 3.1 2.1 1.8 3.1 40.1 25.9 25.97 $\pm$ 0.78
[…]
Table 2: Energy (E) and force (F) root-mean-square errors (RMSEs) for the 3BPA data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 450 configurations randomly drawn from the training data set collected at 300 K, while the remaining 50 are used for early stopping. All values for ICTP models are obtained by averaging over five independent runs. For MACE and NequIP, the results are reported for three independent runs. The standard deviation is provided if it is available. Best performances, considering the standard deviation, are highlighted in bold. Inference time and memory consumption are measured for a batch size of 100. Inference time is reported per structure in ms, while memory consumption is provided for the entire batch in GB.
[…]
300 K E 2.70 $\pm$ 0.22 2.70 $\pm$ 0.08 2.98 $\pm$ 0.34 2.81 $\pm$ 0.18 6.3 3.0 $\pm$ 0.2 3.28 $\pm$ 0.10
F 9.45 $\pm$ 0.29 9.39 $\pm$ 0.31 9.57 $\pm$ 0.20 9.47 $\pm$ 0.42 21.4 8.8 $\pm$ 0.3 10.77 $\pm$ 0.19
600 K E 10.74 $\pm$ 0.31 10.38 $\pm$ 0.80 10.29 $\pm$ 0.90 $1$$\pm$ 1.41 18.0 9.7 $\pm$ 0.5 11.16 $\pm$ 0.14
F 22.99 $\pm$ 0.64 22.87 $\pm$ 0.91 23.03 $\pm$ 0.76 23.27 $\pm$ 1.45 45.2 21.8 $\pm$ 0.6 26.37 $\pm$ 0.09
1200 K E 29.80 $\pm$ 0.92 30.84 $\pm$ 1.87 31.32 $\pm$ 1.80 31.15 $\pm$ 1.58 58.0 29.8 $\pm$ 1.0 38.52 $\pm$ 1.63
F 62.82 $\pm$ 1.23 64.54 $\pm$ 3.88 65.36 $\pm$ 3.47 65.22 $\pm$ 3.52 113.8 62.0 $\pm$ 0.7 76.18 $\pm$ 1.11
Dihedral slices E 9.82 $\pm$ 0.79 10.64 $\pm$ 1.07 13.03 $\pm$ 3.44 8.56 $\pm$ 1.53 7.8 $\pm$ 0.6 23.2 [33]
F 17.52 $\pm$ 0.54 17.18 $\pm$ 0.81 19.31 $\pm$ 0.83 17.69 $\pm$ 1.29 16.5 $\pm$ 1.7 23.1 [33]
Inference time 6.45 $\pm$ 0.50 5.31 $\pm$ 0.02 3.51 $\pm$ 0.22 4.66 $\pm$ 0.05 24.3b 103.5b [33]
Memory consumption 49.66 $\pm$ 0.00 42.01 $\pm$ 0.11 39.08 $\pm$ 0.00 36.26 $\pm$ 0.00
a During inference time measurements with the MACE source code, we were not able to reproduce the original results [33]. Thus, we re-run MACE experiments using the same training setup as for ICTP; see Appendix D.2.
b The original publication did not report the batch size used to measure inference time [33]. Therefore, the values provided are used solely to demonstrate the relative computational cost of MACE and NequIP.
Extrapolation to out-of-domain data. We further assess the performance of ICTP models using the 3BPA data set [48]. The training data set comprises 500 configurations, total energies, and atomic forces acquired from molecular dynamics at 300 K. The test data set is obtained from simulations at 300 K, 600 K, and 1200 K. We also test our models using energies and forces along dihedral rotations of the molecule. Table 2 shows that ICTP models trained using 450 configurations perform on par with state-of-the-art spherical models, similar to the results for rMD17. However, we were not able to reproduce the original results using the current MACE source code and the described training setup [33]. Therefore, for a fair comparison, we unified the training setup for ICTP and MACE (see Appendix D.2) and Table 2 reports the corresponding results for 450 training configurations. In Table A1, we present the results obtained for ICTP and MACE trained using 50 configurations.
[…]
Figure 2: Potential energy profiles for three cuts through the 3BPA molecule’s potential energy surface. All models are trained using 50 configurations, and additional 50 are used for early stopping. The 3BPA molecule, including the three dihedral angles ($\alpha$, $\beta$, and $\gamma$), provided in degrees , is shown as an inset. The color code of the inset molecule is C grey, O red, N blue, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. Shaded areas denote standard deviations across five independent runs.
[…]
Table 3: Energy (E) and force (F) root-mean-square errors (RMSEs) for the acetylacetone data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 450 configurations randomly drawn from the training data set collected at 300 K, while the remaining 50 are used for early stopping. All values for ICTP models are obtained by averaging over five independent runs. For MACE and NequIP, the results are reported for three independent runs. The standard deviation is provided if it is available. Best performances, considering the standard deviation, are highlighted in bold.
[…]
300 K E 0.75 $\pm$ 0.04 0.76 $\pm$ 0.03 0.77 $\pm$ 0.04 0.75 $\pm$ 0.05 0.9 $\pm$ 0.03 0.81 $\pm$ 0.05
F 5.08 $\pm$ 0.11 5.17 $\pm$ 0.10 5.18 $\pm$ 0.16 5.00 $\pm$ 0.17 5.1 $\pm$ 0.10 5.90 $\pm$ 0.46
600 K E 5.39 $\pm$ 1.22 4.43 $\pm$ 0.34 5.12 $\pm$ 0.29 4.96 $\pm$ 0.64 4.6 $\pm$ 0.3 6.04 $\pm$ 1.54
F 23.21 $\pm$ 1.96 22.90 $\pm$ 1.62 24.05 $\pm$ 1.71 23.25 $\pm$ 1.82 22.4 $\pm$ 0.9 27.8 $\pm$ 4.03
Number of parameters 2,774,800 2,736,400 2,648,080 2,803,984 2,803,984 3,190,488
a Similar to Table 2, we re-run MACE experiments using the same training setup as for ICTP; see Appendix D.2.
[…]
6 Conclusions and limitations
This work introduces many-body equivariant MPNNs based on higher-rank irreducible Cartesian tensors, offering an alternative to spherical models. We assess the performance of resulting MPNNs using molecular benchmark data sets, such as rMD17, 3BPA, and acetylacetone. In these experiments, MPNNs based on irreducible Cartesian tensors show a lower computational cost of individual operations compared to spherical counterparts. Furthermore, we demonstrate that these Cartesian models achieve accuracy and generalization capability on par with or better than state-of-the-art spherical models while memory consumption is comparable. Our results hold across the typical range of tensor ranks used in modeling many-body interactions and relevant physical properties, i.e., $L\leq 4$.
Limitations. We emphasize our focus on introducing MPNNs based on irreducible Cartesian tensors and prove their equivariance. We adapted the MACE architecture, which uses only tensors of even parity, to enable a fair comparison with spherical models. Further modifications to the architecture are possible, which we expect to exploit the full potential of irreducible Cartesian tensors. Future work should also extend the assessment of MPNNs based on irreducible Cartesian tensors using larger and more diverse atomic systems.
Code availability
The source code will be made publicly available upon publication.
Acknowledgements
MN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2075 – 390740016 and the Stuttgart Center for Simulation Science (SimTech).
Appendix A Background
Message-passing neural networks. Message-passing neural networks (MPNNs) learn node representations in a graph by iteratively processing local information sent by the nodes’ neighbors. The initial features of node $u$ are represented as the vector $\mathbf{x}_{u}$, and undirected edges $\{u,v\}$ connect pairs of nodes $u,v$. A node $v$ belongs to the neighborhood of node $u$, denoted as $\mathcal{N}(u)$, if there exists an edge $\{u,v\}$ in the graph. Typically, the $(t+1)$-th message-passing layer computes a new node $u$’s representation $\mathbf{h}^{(t+1)}_{u}$ by applying a permutation invariant aggregation function over the neighbors [16, 21]
$\mathbf{h}_{u}^{(t+1)}=\phi^{(t)}\Big{(}\mathbf{h}^{(t)}_{u},\sum\nolimits_{v\in\mathcal{N}(u)}\psi^{(t)}\big{(}\mathbf{h}^{(t)}_{u},\mathbf{h}^{(t)}_{v}\big{)}\Big{)},$
where $\phi^{(t)},\psi^{(t)}$ are often implemented as learnable fully-connected neural networks (NNs), and $\mathbf{h}_{u}^{(0)}=\mathbf{x}_{u}$. To learn a mapping from a learned representation $\mathbf{h}_{u}^{(t)}$ to the atoms’ energies, we can couple $T$ message-passing layers with corresponding readout functions $\mathcal{R}_{t},\,t\in\{1,\dots,T\}$ such that $E_{u}=\sum_{t=1}^{T}\mathcal{R}_{t}\big{(}\mathbf{h}^{(t)}_{u}\big{)}$.
Many-body interatomic potentials. Interatomic potentials approximate the potential energy of atoms—the electronic ground state energy—as a function of their coordinates [77]. Many-body potentials naturally arise because the interaction between two atoms is influenced by the presence of additional atoms, changing their electronic structure. This concept is formalized by expanding the atomic energy $E_{u}$ of a many-atom system into a series of two-body, three-body, and higher-body-order contributions
$E_{u}=E^{(1)}(\mathbf{r}_{u})+\sum_{v_{1}}E^{(2)}(\mathbf{r}_{u},\mathbf{r}_{v_{1}})+\sum_{v_{1}<v_{2}}E^{(3)}(\mathbf{r}_{u},\mathbf{r}_{v_{1}},\mathbf{r}_{v_{2}})+\cdots,$
where $\mathbf{r}_{u}$ represents the position of atom $u$ and the superscript $\nu$ in $E^{(\nu)}$ indicates the order of the many-body interaction. In the absence of external fields, $E^{(\nu)}$ contributions to the atomic energy are invariant to rotations, and the two-body potential depends only on distances $r_{uv}=\lVert\mathbf{r}_{u}-\mathbf{r}_{v}\rVert_{2}$
$E_{u}=\sum_{v_{1}}E^{(2)}(r_{uv_{1}})+\sum_{v_{1}<v_{2}}E^{(3)}(r_{uv_{1}},r_{uv_{2}},r_{v_{1}v_{2}})+\cdots.$
Expansions of this form have found a broad application in constructing machine-learned interatomic potentials (MLIPs), significantly advancing the field [50, 40].
Many-body message passing. Designing accurate and computationally efficient interatomic potential models for interacting many-body systems requires including higher-body-order energy contributions and, thus, higher-body-order learnable features. Recently, a new message construction mechanism has been proposed by expanding the messages $\mathbf{m}_{u}^{(t)}$ to include many-body contributions [33]
$\mathbf{m}_{u}^{(t)}=\sum\limits_{v_{1}}\psi_{2}^{(t)}\big{(}\mathbf{h}^{(t)}_{u},\mathbf{h}^{(t)}_{v_{1}}\big{)}+\sum\limits_{v_{1},v_{2}}\psi_{3}^{(t)}\big{(}\mathbf{h}^{(t)}_{u},\mathbf{h}^{(t)}_{v_{1}},\mathbf{h}^{(t)}_{v_{2}}\big{)}+\cdots+\sum\limits_{v_{1},\cdots,v_{\nu}}\psi_{\nu+1}^{(t)}\big{(}\mathbf{h}^{(t)}_{u},\mathbf{h}^{(t)}_{v_{1}},\cdots,\mathbf{h}^{(t)}_{v_{\nu}}\big{)},$
where $(\nu+1)$ denotes the order of many-body interactions, defining the number of contracted tensors. Using $\sum_{v_{1},\cdots,v_{\nu}}$ instead of $\sum_{v_{1}<\cdots<v_{\nu}}$ circumvents the exponential increase in computational cost with $\nu$. It allows exploiting the product structure of many-body features, which differs from other approaches [69, 68, 70].
Appendix B Methods
B.1 Normalization constants
The following provides the normalization constants $C_{l_{1}l_{2}l_{3}}$ and $D_{l_{1}l_{2}l_{3}}$ for even and odd irreducible Cartesian tensor products in Eqs. (2) and (3), respectively. The respective normalization constants read [45]
$\begin{split}C_{l_{1}l_{2}l_{3}}&=\frac{l_{1}!l_{2}!(2l_{3}-1)!!((L_{1}+1)/2)!((L_{2}+1)/2)!}{l_{3}!L_{1}!!L_{2}!!L_{3}!!(L/2)!},\\ D_{l_{1}l_{2}l_{3}}&=\frac{2l_{1}!l_{2}!(2l_{3}-1)!!(L_{1}/2)!(L_{2}/2)!}{(l_{3}-1)!(L_{1}+1)!!(L_{2}+1)!!(L_{3}+1)!!((L+1)/2)!},\end{split}$
with $L=l_{1}+l_{2}+l_{3}$ and $L_{i}=L-2l_{i}-1$. Here, $C_{l_{1}l_{2}l_{3}}$ is defined such that an $l_{3}$-fold contraction of the tensor $\mathbf{z}_{l_{3}}$, obtained through the irreducible Cartesian tensor product between $\mathbf{x}_{l_{1}}$ and $\mathbf{y}_{l_{2}}$ ($\mathbf{x}_{l_{1}}$ and $\mathbf{y}_{l_{2}}$ are obtained using Eq. (1)), with the unit vector $\hat{\mathbf{r}}$ yields unity. Please refer to the original publication for the motivation behind $D_{l_{1}l_{2}l_{3}}$ [45].
B.2 Further details on the equivariant message passing
The following provides additional details on the employed equivariant message-passing architecture. Learnable weights employed in Eqs. (4), (5), and (6) (i.e., $W_{kk^{\prime}l_{2}}^{(t)}$, $W_{kZ_{v}}^{(t)}$, and $W_{kk^{\prime}l_{\xi}}^{(t)}$ with $k,k^{\prime}$ running over feature channels) are initialized by picking the respective entries from a normal distribution with zero mean and unit variance.
Many-body message-passing. The many-body equivariant features, represented by an irreducible Cartesian tensor of rank $L$ and obtained through the $\nu$-fold tensor product in Eq. (6), are combined using the linear expansion
$\big{(}\mathbf{m}_{ukL}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}}=\Big{(}\sum_{\nu}\sum_{\eta_{\nu}}W_{Z_{u}\eta_{\nu}kL}^{(t)}\mathbf{B}_{u\eta_{\nu}kL}^{(t)}\Big{)}_{i_{1}i_{2}\cdots i_{L}},$ (A1)
where $W_{Z_{u}kL\eta_{\nu}}^{(t)}$ denotes a learnable weight matrix which depends on the chemical element $Z_{u}$ and rank $L$ and which elements are initialized by picking the respective entries from a normal distribution with zero mean and variance of $1/\mathrm{len}(\eta_{\nu})$. The updated node embeddings are further obtained as a linear function of $\big{(}\mathbf{m}_{ukL}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}}$ and the residual connection [79, 33]
$\big{(}\mathbf{h}_{ukL}^{(t+1)}\big{)}_{i_{1}i_{2}\cdots i_{L}}=\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{kk^{\prime}L}^{(t)}\big{(}\mathbf{m}_{uk^{\prime}L}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}}+\frac{1}{\sqrt{d_{t}N_{Z}}}\sum_{k^{\prime}}W_{Z_{u}kk^{\prime}L}^{(t)}\big{(}\mathbf{h}_{uk^{\prime}L}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}},$ (A2)
where $N_{Z}$ denotes the number of atom types and learnable weights $W_{kk^{\prime}L}^{(t)}$ and $W_{Z_{u}kk^{\prime}L}^{(t)}$ are initialized by picking the respective entries from a normal distribution with zero mean and unit variance.
Full and symmetric product basis. In the MACE architecture, the authors pre-compute products between the generalized Clebsch–Gordan coefficients, which define the interaction between $l_{1}$- and $l_{2}$-rank spherical tensors, and the learnable weight $W_{Z_{u}\eta_{\nu}kL}^{(t)}$ of the linear expansion in Eq. (A1[33]. This approach reduces the computational cost of constructing the product basis with spherical tensors as the effective number of evaluated tensor products is smaller. For irreducible Cartesian tensors, operations like matrix-vector and matrix-matrix products define the interaction between $l_{1}$- and $l_{2}$-rank tensors. Thus, an equivalent operation to those proposed for spherical tensors is generally impossible or would lead to an architecture different from MACE.
Note that one of our main goals is to demonstrate that a many-body equivariant message-passing architecture defined using higher-rank irreducible Cartesian tensors can be as expressive as the one using spherical tensors. Therefore, we compute $\mathbf{m}_{ukL}^{(t)}$ by directly evaluating the $\nu$-fold tensor product, which effectively leads to a larger number of operations than for MACE. We refer to the models based on this architecture as irreducible Cartesian tensor potentials (ICTPs) with the full product basis, or ICTPfull. We also noticed that the number of tensor products, $\mathrm{len}(\eta_{\nu})$, leading to the tensor of rank $L$, can be reduced by the symmetry of the two-fold tensor products, i.e., $\mathbf{A}_{l_{1}}\otimes\mathbf{A}_{l_{2}}=\mathbf{A}_{l_{2}}\otimes\mathbf{A}_{l_{1}}$ if $\mathbf{A}_{l_{1}}$ and $\mathbf{A}_{l_{2}}$ are symmetric. We refer to this design choice as ICTPs with the symmetric product basis or ICTPsym.
Coupled and uncoupled feature channels. Using coupled feature channels instead of uncoupled ones in Eq. (A1) can improve the performance of the final model. Additionally, the number of feature channels and, thus, the overall computational cost can be reduced when constructing the product basis. However, the number of parameters and, thus, the expressiveness of the model can be preserved. Specifically, we can define a linear expansion for combining many-body equivariant features as
$\big{(}\mathbf{m}_{ukL}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{L}}=\Big{(}\sum_{\nu}\sum_{\eta_{\nu}}\sum_{k^{\prime}}W_{Z_{u}\eta_{\nu}kk^{\prime}L}^{(t)}\mathbf{B}_{u\eta_{\nu}k^{\prime}L}^{(t)}\Big{)}_{i_{1}i_{2}\cdots i_{L}},$ (A3)
where $W_{Z_{u}kk^{\prime}L\eta_{\nu}}^{(t)}$ denotes a learnable weight matrix which depends on the chemical element $Z_{u}$ and rank $L$ and which elements are initialized by picking the respective entries from a normal distribution with zero mean and variance of $1/(\sqrt{d_{t}}\times\mathrm{len}(\eta_{\nu}))$. The construction of the product basis in the latent feature space requires encoding the two-body features $\mathbf{A}_{ukl_{\xi}}^{(t)}$ using the learnable weight matrices in Eq. (6). The linear expansion is then constructed in the latent feature space and decoded using the learnable weight matrices of the linear function in Eq. (A2). We refer to this design choice as ICTPs with the product basis constructed in the latent feature space or ICTPlt. Combining it with the symmetric product basis, we obtain ICTPsym+lt.
Readout. Atomic energies expanded into a series of many-body contributions, $E_{u}=E_{u}^{(0)}+E_{u}^{(1)}+\cdots+E_{u}^{(T)}$, are obtained by applying readout functions $\mathcal{R}_{t}$ to node features with $L=0$, which are invariant to rotations
$E_{u}^{(t)}=\mathcal{R}_{t}\Big{(}\big{\{}\mathbf{h}_{uk(L=0)}^{(t)}\big{\}}_{k}\Big{)}=\begin{cases}\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{k^{\prime}}^{(t)}h_{uk^{\prime}(L=0)}^{(t)}&\text{if }t<T,\\ \mathrm{NN}^{(t)}\Big{(}\big{\{}\mathbf{h}_{uk(L=0)}^{(t)}\big{\}}_{k}\Big{)}&\text{if }t=T,\end{cases}$ (A4)
with learnable weights $W_{k^{\prime}}^{(t)}$ or those of $\mathrm{NN}^{(t)}$ initialized by picking the respective entries from a normal distribution with zero mean and unit variance. Linear readout functions for $t<T$ preserve the many-body orders in $\mathbf{h}_{uk(L=0)}^{(t)}$, while a one-layer fully-connected NN is used for the last message-passing layer and accounts for the residual higher-order terms in the expansion [32].
B.3 Asymptotic computational complexity
We analyze the computational complexity of the irreducible Cartesian tensor product with respect to the maximum rank $L$ used. For each tuple of $\left(i_{1},\cdots,i_{L}\right)$ of a rank-$L$ tensor in Eq. (2), the single contraction term $\left(\mathbf{x}_{L}\cdot(k+m)\cdot\mathbf{y}_{L}\right)$ has a computational complexity of $\mathcal{O}\left(3^{k+m}\right)$. The total computational complexity is $\mathcal{O}\left(3^{L}\right)$ for even products and $\mathcal{O}\left(3^{L+1}\right)$ for odd ones—due to the additional double contraction with the Levi-Civita symbol. The computation of the set of permutations over the $L$ unsymmetrized indices scales as $L!/\left(2^{L/2}\left(L/2\right)!\right)$ for even products and $L!/\left(2^{\left(L-1\right)/2}\left(\left(L-1\right)/2\right)!\right)$ for odd ones [45]. Because each final rank-$L$ tensor has $3^{L}$ elements, the total complexity for computing an irreducible Cartesian tensor of rank $L$ is $\mathcal{O}\left(9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$.
Recent state-of-the-art equivariant message-passing architectures employ tensors of rank $L\leq 4$ [33, 32, 78]. Also, physical properties are represented by tensors of rank $L\leq 4$. For these spherical tensors, the number of calculations required to obtain an element of a new $L$-rank tensor is $\left(2L+1\right)^{5}$, or $\left(2L+1\right)^{6}$ for a $\left(2L+1\right)$-dimensional object $\mathbf{Y}^{L}$. Thus, the irreducible Cartesian tensor product is expected to be more computationally efficient for $L\leq 4$, assuming similar multiplicative factors and negligible sub-leading terms. Compared to the Gaunt tensor product with $(2L+1)^{3}$, irreducible Cartesian tensors are expected to be less computationally efficient for $L\geq 3$. However, while the Gaunt tensor product is restricted only to even parity, i.e., it excludes tensors with odd $l_{1}+l_{2}-l_{3}=2k+1$ (e.g., pseudo-vectors), the irreducible Cartesian tensor product naturally allows their computation.
As regards the cost of performing message passing at each layer, the general neighbors’ aggregation of Eq. (4) has a cost of $\mathcal{O}\left(\mathcal{E}9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$, where $\mathcal{E}$ is the number of edges in the system. Similarly, computing the many-body features tensor via Eq. (6) requires to perform $\nu$ Cartesian tensor products for all $\mathcal{K}=\mathrm{len}(\eta_{\nu})$ possible $\nu$-fold tensor products, resulting in an asymptotic complexity of $\mathcal{O}\left(\mathcal{M}\mathcal{K}\nu 9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$ with $\mathcal{M}$ denoting the number of nodes. Thus, the computational complexity for one ICTP message-passing layer is $\mathcal{O}\left(\left(\mathcal{E}+\mathcal{M}\mathcal{K}\nu\right)9^{L}L!/\left(2^{L/2}\left(L/2\right)!\right)\right)$.
Finally, memory consumption for Cartesian tensors is often believed to be less advantageous than that of spherical tensors. However, for contracting two rank-$L$ spherical tensors, space complexity is about $(2L+1)^{2}$, resulting from the definition of the Clebsch–Gordan tensor product in Section 2, i.e., due to the computation of intermediate tensors $\mathbf{Y}^{L}\otimes\mathbf{Y}^{L}\in\mathbb{R}^{(2L+1)\times(2L+1)}$. For a Cartesian tensor, the space complexity is about $3^{L}$. Thus, the irreducible Cartesian tensor product can be expected to require the same or less memory for $L\leq 4$ than the Clebsch–Gordan counterpart.
Appendix C Proof of Cartesian message-passing equivariance to actions of the orthogonal group
We first recap the standard action of the $\mathrm{O}(3)$ group onto $\left(\mathbb{R}^{3}\right)^{\otimes l}$. Let $D_{\mathcal{X}}$ be a representation of the orthogonal group $\mathrm{O}(3)$ (also in line with Section 2), i.e.,
$D_{\mathcal{X}}:\mathrm{O}(3)\rightarrow\mathbb{R}^{3\times 3},\ \ \ g\mapsto D_{\mathcal{X}}[g]=R.$
We define the action of the $\mathrm{O}(3)$ group onto $\left(\mathbb{R}^{3}\right)^{\otimes l}$ as follows
$\phi:\mathrm{O}(3)\times\left(\mathbb{R}^{3}\right)^{\otimes l}\rightarrow\left(\mathbb{R}^{3}\right)^{\otimes l},\ \ \ \phi\left(g,\mathbf{T}_{l}\right)_{i_{1}i_{2}\cdots i_{l}}\coloneqq\sum_{j_{1}}\cdots\sum_{j_{l}}R_{i_{1}j_{1}}\cdots R_{i_{l}j_{l}}\left(\mathbf{T}_{l}\right)_{j_{1}\cdots j_{l}},$
where $\mathbf{T}_{l}\in\left(\mathbb{R}^{3}\right)^{\otimes l}$. We hereinafter denote $\phi\left(g,\mathbf{T}_{l}\right)$ also by $R\mathbf{T}_{l}$. The outer product $\otimes$ is equivariant to this action, and $R$ acts trivially on the $3\times 3$-identity matrix.
Lemma C.1.
Let $R$ be a representation of an element of the orthogonal group $\mathrm{O}(3)$. Then,
  • (i)
    $\forall\,\mathbf{T}_{l_{1}}\in\left(\mathbb{R}^{3}\right)^{\otimes l_{1}}$ and $\forall\,\mathbf{T}_{l_{2}}\in\left(\mathbb{R}^{3}\right)^{\otimes l_{2}}$
    $R\left(\mathbf{T}_{l_{1}}\otimes\mathbf{T}_{l_{2}}\right)=\left(R\mathbf{T}_{l_{1}}\right)\otimes\left(R\mathbf{T}_{l_{2}}\right).$
  • (ii)
    For the $3\times 3$-identity matrix $\mathbf{I}$, we have
    $R\mathbf{I}=\mathbf{I}.$
Proof.
(i) We first show that $R\left(\mathbf{T}_{l_{1}}\otimes\mathbf{T}_{l_{2}}\right)=\left(R\mathbf{T}_{l_{1}}\right)\otimes\left(R\mathbf{T}_{l_{2}}\right)$
$\begin{split}&\left(\left(R\mathbf{T}_{l_{1}}\right)\otimes\left(R\mathbf{T}_{l_{2}}\right)\right)_{i_{1}\cdots i_{l_{1}}i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}=\\ &\left(R\mathbf{T}_{l_{1}}\right)_{i_{1}\cdots i_{l_{1}}}\left(R\mathbf{T}_{l_{2}}\right)_{i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}=\\ &\left(\sum_{j_{1}}\cdots\sum_{j_{l_{1}}}R_{i_{1}j_{1}}\cdots R_{i_{l}j_{l_{1}}}\left(\mathbf{T}_{l_{1}}\right)_{j_{1}\cdots j_{l_{1}}}\right)\left(\sum_{j_{l_{1}+1}}\cdots\sum_{j_{l_{1}+l_{2}}}R_{i_{l_{1}+1}j_{l_{1}+1}}\cdots R_{i_{l_{1}+l_{2}}j_{l_{1}+l_{2}}}\left(\mathbf{T}_{l_{2}}\right)_{j_{l_{1}+1}\cdots j_{l_{1}+l_{2}}}\right)=\\ &\sum_{j_{1}}\cdots\sum_{j_{l_{1}}}\sum_{j_{l_{1}+1}}\cdots\sum_{j_{l_{1}+l_{2}}}R_{i_{1}j_{1}}\cdots R_{i_{l_{1}}j_{l_{1}}}R_{i_{l_{1}+1}j_{l_{1}+1}}\cdots R_{i_{l_{1}+l_{2}}j_{l_{1}+l_{2}}}\left(\mathbf{T}_{l_{1}}\right)_{j_{1}\cdots j_{l_{1}}}\left(\mathbf{T}_{l_{2}}\right)_{j_{l_{1}+1}\cdots j_{l_{1}+l_{2}}}\\ &=\left(R\left(\mathbf{T}_{l_{1}}\otimes\mathbf{T}_{l_{2}}\right)\right)_{i_{1}\cdots i_{l_{1}}i_{l_{1}+1}\cdots i_{l_{1}+l_{2}}}.\end{split}$
(ii) With $\sum_{i}R_{ij}R_{ik}=\delta_{jk}$, or equivalently $R^{T}R=\mathbf{I}$, we get
$(R\mathbf{I})_{i_{1}i_{2}}=\sum_{j_{1},j_{2}}R_{i_{1}j_{1}}R_{i_{2}j_{2}}\delta_{j_{1}j_{2}}=\sum_{j_{1}}R_{i_{1}j_{1}}R_{i_{2}j_{1}}=\delta_{i_{1}i_{2}}.$
Using Lemma C.1, we can show that the irreducible Cartesian tensor operator $\mathbf{T}_{l}:\mathbb{R}^{3}\rightarrow(\mathbb{R}^{3})^{\otimes l}$ is equivariant to actions of the $\mathrm{O}(3)$ group.
Proposition C.2.
Let $l\geq 0$ be a positive integer. Then, the irreducible Cartesian tensor operator $\mathbf{T}_{l}:\mathbb{R}^{3}\rightarrow(\mathbb{R}^{3})^{\otimes l}$ is equivariant to actions of the $\mathrm{O}(3)$ group.
Proof.
It suffices to show that the map $\hat{\mathbf{r}}\mapsto\hat{\mathbf{r}}^{\otimes(l-2m)}\otimes\mathbf{I}^{\otimes m}$ is equivariant to actions of the $\mathrm{O}(3)$ group $\forall\,l\geq 1$ and $\forall\,m\leq\lfloor l/2\rfloor$ ($m\geq 0$).
[…]
Our next claim is that for $l_{3}\in\left\{\lvert l_{1}-l_{2}\rvert,\lvert l_{1}-l_{2}\rvert+1,\cdots,l_{1}+l_{2}\right\}$ the irreducible Cartesian tensor product $\otimes_{\mathrm{Cart}}$ defined in Eqs. (2) and (3) is equivariant to actions of the $\mathrm{O}(3)$ group, i.e., the following diagram commutes
[…]
The proof for this claim is very similar to Proposition C.2. The only difference is to show the equivariance for the $(k+m)$-fold tensor contraction.
Proposition C.3.
The irreducible Cartesian tensor product $\otimes_{\mathrm{Cart}}:\left(\mathbb{R}^{3}\right)^{\otimes l_{1}}\times\left(\mathbb{R}^{3}\right)^{\otimes l_{2}}\rightarrow\left(\mathbb{R}^{3}\right)^{\otimes l_{3}}$ makes the above diagram commute.
Proof.
It suffices to show that the $(k+m)$-fold tensor contraction is equivariant to actions of the $\mathrm{O}(3)$ group. For $\mathbf{x}_{l_{1}}\in(\mathbb{R}^{3})^{\otimes l_{1}}$, $\mathbf{y}_{l_{2}}\in(\mathbb{R}^{3})^{\otimes l_{2}}$, and $R$ being a representation of an element of the orthogonal group $\mathrm{O}(3)$, we can write
$\begin{split}&\left(\left(R\mathbf{x}_{l_{1}}\right)\cdot(s)\cdot\left(R\mathbf{y}_{l_{2}}\right)\right)_{\beta_{1}\cdots\beta_{l_{1}-s}\delta_{1}\cdots\delta_{l_{2}-s}}=\\ &\sum_{\alpha_{1},\cdots,\alpha_{s}}\left(R\mathbf{x}_{l_{1}}\right)_{\alpha_{1}\cdots\alpha_{s}\beta_{1}\cdots\beta_{l_{1}-s}}\left(R\mathbf{y}_{l_{2}}\right)_{\alpha_{1}\cdots\alpha_{s}\delta_{1}\cdots\delta_{l_{2}-s}}=\\ &\sum_{\alpha_{1},\cdots,\alpha_{s}}\left(\sum_{\begin{subarray}{c}\gamma_{1},\cdots,\gamma_{s}\\ \eta_{1},\cdots,\eta_{l_{1}-s}\end{subarray}}R_{\alpha_{1}\gamma_{1}}\cdots R_{\alpha_{s}\gamma_{s}}R_{\beta_{1}\eta_{1}}\cdots R_{\beta_{l_{1}-s}\eta_{l_{1}-s}}\left(\mathbf{x}_{l_{1}}\right)_{\gamma_{1}\cdots\gamma_{s}\eta_{1}\cdots\eta_{l_{1}-s}}\right)\times\\ &\hskip 34.1433pt\left(\sum_{\begin{subarray}{c}\tilde{\gamma}_{1},\cdots,\widetilde{\gamma}_{s}\\ \widetilde{\eta}_{1},\cdots,\widetilde{\eta}_{l_{2}-s}\end{subarray}}R_{\alpha_{1}\widetilde{\gamma}_{1}}\cdots R_{\alpha_{s}\widetilde{\gamma}_{s}}R_{\delta_{1}\widetilde{\eta}_{1}}\cdots R_{\delta_{l_{2}-s}\widetilde{\eta}_{l_{2}-s}}\left(\mathbf{y}_{l_{2}}\right)_{\widetilde{\gamma}_{1}\cdots\widetilde{\gamma}_{s}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}}\right)=\\ &\left(\sum_{\begin{subarray}{c}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{1}-s}\end{subarray}}R_{\beta_{1}\eta_{1}}\cdots R_{\beta_{l_{1}-s}\eta_{l_{1}-s}}\right)\left(\sum_{\begin{subarray}{c}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}\end{subarray}}R_{\delta_{1}\widetilde{\eta}_{1}}\cdots R_{\delta_{l_{2}-s}\widetilde{\eta}_{l_{2}-s}}\right)\times\\ &\hskip 135.15059pt\sum_{\begin{subarray}{c}\gamma_{1}\cdots\gamma_{s}\end{subarray}}\left(\mathbf{x}_{l_{1}}\right)_{\gamma_{1}\cdots\gamma_{s}\eta_{1}\cdots\eta_{l_{1}-s}}\left(\mathbf{y}_{l_{2}}\right)_{\gamma_{1}\cdots\gamma_{s}\widetilde{\eta}_{1}\cdots\widetilde{\eta}_{l_{2}-s}}=\\ &\left(R\left(\mathbf{x}_{l_{1}}\cdot(s)\cdot\mathbf{y}_{l_{2}}\right)\right)_{\beta_{1}\cdots\beta_{l_{1}-s}\delta_{1}\cdots\delta_{l_{2}-s}}.\end{split}$
A similar derivation applies to the odd case (3), which completes the proof. ∎
We finally give a proof for Proposition 4.1, i.e., the equivariance of message-passing layers based on irreducible Cartesian tensors to actions of the $\mathrm{O}(3)$ group.
Proof of Proposition 4.1.
Since a message-passing layer in Section 4.2 and Appendix B.2 is defined by stacking layers corresponding to Eqs. (4), (5), (6), (A1), (A2), and (A3), we show the equivariance for each of these equations. We show the equivariance by induction.
  • Case $t=1$:
    • Equivariance of Eq. (5): Since learnable weights $W_{kZ_{v}}$ and learnable (invariant) radial basis $R_{kl_{1}}^{(1)}(r_{uv})$ are defined for each feature channel $k$, the weights are multiplied with $\big{(}\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}$ as scalars. Therefore,
      $\big{(}\mathbf{A}_{ukl_{1}}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}=\sum_{v\in\mathcal{N}(u)}R_{kl_{1}}^{(1)}(r_{uv})\big{(}\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\big{)}_{i_{1}i_{2}\cdots i_{l_{1}}}W_{kZ_{v}}$
      is equivariant to actions of the $\mathrm{O}(3)$ group.
    • Equivariance of Eq. (6): The proof is similar to the case of Eq. (5), since we again have learnable parameter $W_{kk^{\prime}l_{\nu}}^{(1)}$ for each $\mathbf{A}_{uk^{\prime}l_{\nu}}^{(1)}$. Therefore, noting that the Cartesian tensor product is equivariant to actions of the $\mathrm{O}(3)$ group, the following function
      $\big{(}\mathbf{B}_{u\eta_{\nu}kL}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{L}}=\big{(}\tilde{\mathbf{A}}_{ukl_{1}}^{(1)}\otimes_{\mathrm{Cart}}\cdots\otimes_{\mathrm{Cart}}\tilde{\mathbf{A}}_{ukl_{\nu}}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{L}}$
      is also equivariant to actions of the $\mathrm{O}(3)$ group.
    • Equivariance of Eq. (A1): The equation defined by
      $\big{(}\mathbf{m}_{ukL}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{L}}=\Big{(}\sum_{\nu}\sum_{\eta_{\nu}}W_{Z_{u}\eta_{\nu}kL}^{(1)}\mathbf{B}_{u\eta_{\nu}kL}^{(1)}\Big{)}_{i_{1}i_{2}\cdots i_{L}},$
      is equivariant to actions of the $\mathrm{O}(3)$ group, since for each set of indices $u$, $\eta_{\nu}$, $k$, and $L$ we have a scalar learnable parameter $W_{Z_{u}\eta_{\nu}kL}^{(1)}$.
    • Equivariance of Eq. (A2): The respective parameters $W_{kk^{\prime}L}^{(1)}$ and $W_{Z_{u}kk^{\prime}L}^{(1)}$ in Eq. (A2)
      $\big{(}\mathbf{h}_{ukL}^{(2)}\big{)}_{i_{1}i_{2}\cdots i_{L}}=\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{kk^{\prime}L}^{(1)}\big{(}\mathbf{m}_{uk^{\prime}L}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{L}}+\frac{1}{\sqrt{d_{t}N_{Z}}}\sum_{k^{\prime}}W_{Z_{u}kk^{\prime}L}^{(1)}\big{(}\mathbf{h}_{uk^{\prime}L}^{(1)}\big{)}_{i_{1}i_{2}\cdots i_{L}}.$
      are multiplied as scalar with $\mathbf{h}_{uk^{\prime}L}^{(1)}$ and $\mathbf{m}_{uk^{\prime}L}^{(1)}$, which are shown to be equivariant to actions of the $\mathrm{O}(3)$ group. Thus, $\mathbf{h}_{ukL}^{(2)}$ is equivariant to actions of the $\mathrm{O}(3)$ group as a function of $\mathbf{h}_{uk^{\prime}L}^{(1)}$ and $\mathbf{m}_{uk^{\prime}L}^{(1)}$.
    • Equivariance of Eq. (A3): The equivariance of Eq. (A3) to actions of the $\mathrm{O}(3)$ group is also immediate since learnable parameters apply to each tuple of $(i_{1},i_{2},\cdots,i_{L})$ as a scalar.
  • General case $t>1$: The equivariance of Eqs. (4), (5), (6), (A1), (A2), and (A3) to actions of the $\mathrm{O}(3)$ group for arbitrary $t>1$ follows in a similar way to those obtained for the $t=1$ case. However, we need to show the equivariance of Eq. (4). Suppose that Eqs. (5), (6), (A1), (A2), and (A3) are equivariant to actions of the $\mathrm{O}(3)$ group for $t>1$. Then, for the $(t+1)$-th message-passing layer,
    $\big{(}\mathbf{A}_{ukl_{3}}^{(t)}\big{)}_{i_{1}i_{2}\cdots i_{l_{3}}}=\sum_{v\in\mathcal{N}(u)}\Big{(}R_{kl_{1}l_{2}l_{3}}^{(t)}(r_{uv})\mathbf{T}_{l_{1}}(\hat{\mathbf{r}}_{uv})\otimes_{\mathrm{Cart}}\frac{1}{\sqrt{d_{t}}}\sum_{k^{\prime}}W_{kk^{\prime}l_{2}}^{(t)}\mathbf{h}_{vk^{\prime}l_{2}}^{(t)}\Big{)}_{i_{1}i_{2}\cdots i_{l_{3}}},$
    is equivariant to actions of the $\mathrm{O}(3)$ group. Here, the irreducible Cartesian tensor product $\otimes_{\mathrm{Cart}}$ and $\mathbf{T}_{l_{1}}$ are equivariant to actions of the $\mathrm{O}(3)$ group, $\mathbf{h}_{vk^{\prime}l_{2}}^{(t)}$ is equivariant by the assumption, and $W_{kk^{\prime}l_{2}}^{(t)}$ applies to $\mathbf{h}_{vk^{\prime}l_{2}}^{(t)}$ as a scalar.
Appendix D Experiments and results
D.1 Description of the data sets
rMD17 data set. The revised MD17 (rMD17) data set is a collection of structures, energies, and atomic forces of ten small organic molecules obtained from ab initio molecular dynamics (AIMD) [47]. These molecules are derived from the original MD17 data set [47, 80, 81, 82], with 100,000 structures sampled for each. Our models are trained using 950 and 50 configurations for each molecule randomly sampled from the original data set using five random seeds, with 50 additional configurations randomly sampled for early stopping. We use the remaining configurations to test the final models. Table 1 reports the mean absolute errors (MAEs) in total energies and atomic forces averaged over five independent runs, including the standard deviation between them.
3BPA data set. The 3BPA data set comprises structures, energies, and atomic forces of a flexible drug-like organic molecule obtained from AIMD at various temperatures [48]. The training data set consists of 500 configurations sampled at 300 K, while three separate test data sets are obtained from AIMD simulations at 300 K, 600 K, and 1200 K. An additional test data set provides energy values along dihedral rotations of the molecule. This test directly assesses the smoothness and accuracy of the potential energy surface, influencing properties such as binding free energies to protein targets. Our models are trained using 450 and 50 configurations randomly sampled from the training data set using five random seeds, with further 50 configurations reserved for early stopping. Table 2 reports the root-mean-square errors (RMSEs) in total energies and atomic forces averaged over five independent runs, including the standard deviation between them.
Acetylacetone data set. The acetylacetone data set includes structures, energies, and atomic forces of a small reactive molecule obtained from AIMD at various temperatures [32]. The training data set comprises configurations sampled at 300 K, while the test data sets are sampled at 300 K and 600 K. The generalization ability of final models is evaluated using an elevated temperature of 600 K and along two internal coordinates of the molecule: The hydrogen transfer path and a partially conjugated double bond rotation featuring a high rotation barrier. Our models are trained using 450 and 50 configurations randomly sampled from the training dataset using five random seeds, with further 50 configurations reserved for early stopping. Table 3 reports the root-mean-square errors (RMSEs) in total energies and atomic forces averaged over five independent runs, including the standard deviation between them.
D.2 Training details
All ICTP and MACE models employed in this work are trained on a single NVIDIA A100 GPU with 80 GB of RAM. Training times for ICTP and MACE models are typically between 30 minutes and 15 hours, depending on the data set, the data set size, and the employed precision. We use double precision for 3BPA and AcAc data sets and single precision for rMD17, in line with the original experiments [33]. For all experiments, we use two message-passing layers and irreducible Cartesian tensors or spherical tensors of a maximal rank of $l_{\mathrm{max}}=3$ to embed the directional information of atomic distance vectors. For ICTP models with the full and symmetric product basis (ICTPfull and ICTPsym, respectively) and MACE, we employ 256 uncoupled feature channels. For ICTP models with the symmetric product basis evaluated in the latent feature space (ICTPsym+lt), we use 64 coupled feature channels for the Cartesian product basis and 256 for two-body features. Radial features are derived from eight Bessel basis functions with polynomial envelope for the cutoff with $p=5$ [70]. These features are fed into a fully connected NN of size $[64,64,64]$. We apply SiLU non-linearities to the outputs of the hidden layers [83, 84]. The readout function of the first message-passing layer is implemented as a linear layer. The readout function of the second layer is a single-layer fully connected NN with 16 hidden neurons. A cutoff radius of 5 Å is used across all data sets.
All parameters of ICTP and MACE models are optimized by minimizing the combined squared loss on training data $\mathcal{D}_{\mathrm{train}}=\left(\mathcal{X}_{\mathrm{train}},\mathcal{Y}_{\mathrm{train}}\right)$, where $\mathcal{X}_{\mathrm{train}}=\{\mathcal{S}^{(k)}\}_{k=1}^{N_{\mathrm{train}}}$ and $\mathcal{Y}_{\mathrm{train}}=\{E_{k}^{\mathrm{ref}},\{\mathbf{F}_{u,k}^{\mathrm{ref}}\}_{u=1}^{N_{\mathrm{at}}}\}_{k=1}^{N_{\mathrm{train}}}$:
$\mathcal{L}\left(\boldsymbol{\theta},\mathcal{D}_{\mathrm{train}}\right)=\sum_{k=1}^{N_{\mathrm{train}}}\Bigg{[}C_{\mathrm{e}}\Big{\lVert}E_{k}^{\mathrm{ref}}-E(\mathcal{S}^{(k)},\boldsymbol{\theta})\Big{\rVert}_{2}^{2}+C_{\mathrm{f}}\sum_{u=1}^{N_{\mathrm{at}}^{(k)}}\Big{\lVert}\mathbf{F}_{u,k}^{\mathrm{ref}}-\mathbf{F}_{u}\left(\mathcal{S}^{(k)},\boldsymbol{\theta}\right)\Big{\rVert}_{2}^{2}\Bigg{]}.$ (A5)
When training ICTP models, we set $C_{\mathrm{e}}=1/N_{\mathrm{at}}^{(k)}$ and $C_{\mathrm{f}}=10~{}\text{\AA}^{2}$ to balance the relative contributions of total energies and atomic forces, respectively. For MACE, $C_{\mathrm{e}}=1/(B\times N_{\mathrm{at}}^{(k)})$ and $C_{\mathrm{f}}=1000/(B\times 3\times N_{\mathrm{at}}^{(k)})~{}\text{\AA}^{2}$ are used with $B$ denoting the batch size. Here, $E(\mathcal{S}^{(k)},\boldsymbol{\theta})$ and $\mathbf{F}_{u}\left(\mathcal{S}^{(k)},\boldsymbol{\theta}\right)$ are total energies and atomic forces predicted by ICTP or MACE.
All models are trained for 2000 epochs using the AMSGrad variant of Adam [85], with default parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\varepsilon=10^{-8}$. We use a learning rate of 0.01 and a batch size of 5 for all experiments. For evaluations on the validation and test data sets, we use a batch size of 10. The learning rate is reduced using an on-plateau scheduler based on the validation loss with a patience of 50 and a decay factor of 0.8. We utilize an exponential moving average with a weight of 0.99 for evaluation on the validation set and for the final model. Additionally, in line with MACE [33], we apply exponential weight decay of $5\times 10^{-7}$ on the weights of Eqs. (6), (A1), and (A3). Furthermore, we incorporate a per-atom shift of total energies via the average per-atom energy over all the training configurations, including the energies of individual atoms for 3BPA and acetylacetone datasets. If no atomic energies are provided, as for rMD17, the per-atom shift is obtained by solving a linear regression problem [56]. Additionally, a per-atom scale is determined as the root-mean-square of the components of the forces over the training configurations.
D.3 Additional results
Extrapolation to out-of-domain data. Table A1 demonstrates total energy and atomic force RMSEs obtained for ICTP and MACE models trained using 50 configurations randomly drawn from the original data set. ICTP and MACE models perform similarly, considering the standard deviation obtained across five independent runs. However, ICTP models often have lower mean RMSE values compared to MACE.
Figure A1 compares potential energy profiles obtained with ICTP and MACE models trained using 450 configurations. Potential energy cuts at $\beta=120^{\circ}$ and $\beta=180^{\circ}$ are easier tasks for MLIPs, as there are training points in the data set with similar combinations of dihedral angles [32]. In contrast, the potential energy cut at $\beta=150^{\circ}$ is more challenging, with no training points close to it. Notably, Fig. A1 shows that all models produce smooth potential energy profiles close to the reference ones (DFT) for all values of $\beta$. These results again demonstrate excellent extrapolation capabilities of irreducible Cartesian models that are on par with the spherical MACE model.
Table A1: Energy (E) and force (F) root-mean-square errors (RMSEs) for the 3BPA data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 50 molecules randomly drawn from the training data set collected at 300 K, while additional 50 are used for early stopping. All values for ICTP and MACE models are obtained by averaging over five independent runs, with the standard deviation provided if available. Best performances, considering the standard deviation, are highlighted in bold.
[…]
300 K E 14.98 $\pm$ 1.62 13.43 $\pm$ 1.00 16.03 $\pm$ 1.26 14.54 $\pm$ 1.02
F 37.21 $\pm$ 2.14 37.27 $\pm$ 1.63 38.38 $\pm$ 1.55 37.68 $\pm$ 1.33
600 K E 31.68 $\pm$ 3.56 31.63 $\pm$ 3.94 30.74 $\pm$ 1.82 30.71 $\pm$ 3.43
F 69.87 $\pm$ 3.04 68.87 $\pm$ 2.94 69.62 $\pm$ 2.88 69.88 $\pm$ 3.88
1200 K E 92.16 $\pm$ 9.33 86.0 $\pm$ 12.03 78.51 $\pm$ 8.88 83.99 $\pm$ 8.20
F 157.72 $\pm$ 6.53 153.16 $\pm$ 9.62 151.37 $\pm$ 9.59 154.46 $\pm$ 10.84
Dihedral slices E 33.69 $\pm$ 8.03 31.06 $\pm$ 6.27 33.44 $\pm$ 5.56 27.79 $\pm$ 7.41
F 47.12 $\pm$ 5.27 47.68 $\pm$ 2.21 49.08 $\pm$ 4.05 48.91 $\pm$ 4.71
Figure A1: Potential energy profiles for three cuts through the 3BPA molecule’s potential energy surface. All models are trained using 450 configurations, and the remaining 50 are used for early stopping. The 3BPA molecule, including the three dihedral angles ($\alpha$, $\beta$, and $\gamma$), provided in degrees , is shown as an inset. The color code of the inset molecule is C grey, O red, N blue, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. Shaded areas denote standard deviations across five independent runs.
Flexibility and reactivity. Table A2 demonstrates total energy and atomic force RMSEs obtained for ICTP and MACE models trained using 50 configurations randomly drawn from the original data set. Similar to the 3BPA data set, ICTP and MACE models demonstrate comparable accuracy in predicted energies and forces, considering the standard deviation obtained across five independent runs.
Figure A2 further investigates the generalization capabilities of ICTP models trained using 450 configurations, demonstrating potential energy profiles for the rotation around the C-C bond, i.e., the C-C-C-O dihedral angle ($\alpha$), and for the hydrogen transfer. The training data set encompasses dihedral angles less than $30^{\circ}$. Furthermore, the energy barrier of 1 eV is outside the energy range of the training data set obtained at 300 K. As for the hydrogen transfer, the training data does not contain transition geometries, but the reaction still occurs in a region that is not too far from the training data. Overall, ICTP models perform on par with MACE for predicting potential energy profiles for the rotation around the corresponding C-C bond and for the hydrogen transfer.
[…]
Table A2: Energy (E) and force (F) root-mean-square errors (RMSEs) for the acetylacetone data set. E- and F-RMSE are provided in meV and meV/Å, respectively. Results are presented for models trained using 50 configurations randomly drawn from the training data set collected at 300 K, while additional 50 are used for early stopping. All values for ICTP and MACE models are obtained by averaging over five independent runs. Best performances, considering the standard deviation, are highlighted in bold.
[…]
300 K E 4.42 $\pm$ 0.39 4.45 $\pm$ 0.36 4.38 $\pm$ 0.24 4.22 $\pm$ 0.52
F 28.85 $\pm$ 3.00 28.28 $\pm$ 1.45 29.17 $\pm$ 1.65 28.38 $\pm$ 2.74
600 K E 17.61 $\pm$ 3.14 16.13 $\pm$ 1.37 17.3 $\pm$ 2.05 17.53 $\pm$ 3.58
F 75.09 $\pm$ 8.70 69.99 $\pm$ 5.12 74.96 $\pm$ 4.68 74.93 $\pm$ 9.91
Figure A2: Potential energy profiles of (a) the dihedral angle describing the rotation around the C-C bond and (b) hydrogen transfer between two oxygen atoms. All models are trained using 450 molecules, and the remaining 50 are used for early stopping. The acetylacetone molecule, including the dihedral angle in degrees describing the rotation around the C-C bond ($\alpha$), is shown as an inset in (a). The color code of the inset molecule is C grey, O red, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. The histograms demonstrate the distribution of dihedral angles and O-H distances in the training data. Shaded areas denote standard deviations across five independent runs.
Figure A3: Potential energy profiles of (a) the dihedral angle describing the rotation around the C-C bond and (b) hydrogen transfer between two oxygen atoms. All models are trained using 50 molecules, and additional 50 are used for early stopping. The acetylacetone molecule, including the dihedral angle in degrees describing the rotation around the C-C bond ($\alpha$), is shown as an inset in (a). The color code of the inset molecule is C grey, O red, and H white. The reference potential energy profile (DFT) is shown in black. Each profile is shifted such that each model’s lowest energy is zero. The histograms demonstrate the distribution of dihedral angles and O-H distances in the training data set for one random seed used to split the original data. Shaded areas denote standard deviations across five independent runs.
Appendix E Broader social impact
This section discusses the broader social impact of the presented work. Our work has important implications for the chemical sciences and engineering, as many problems in these fields require atomistic simulations; we also discuss it in Section 1. Although this work focuses on standard benchmark data sets, our experiments demonstrate the scalability of our method to larger atomic systems. Beyond constructing machine-learned interatomic potentials, equivariant models based on irreducible Cartesian tensors can be applied for molecular property prediction, protein structure prediction, protein generation, RNA structure ranking, and many more.
Our work has no obvious negative social impact. As long as it is applied to the chemical sciences and engineering in a way that benefits society, it will have positive effects.